OpenStudy (anonymous):

The probability of having a winning raffle ticket is 20%. If you bought 50 tickets, how many winning tickets should you expect to have A.3 tickets B.8 tickets C.10 tickets D.20 tickets

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

anyone plz help

OpenStudy (tkhunny):

What number is 20% of 50? Stop tagging people. Start showing YOUR work.

OpenStudy (anonymous):

sheesh rude

OpenStudy (welshfella):

20% = 20/100 multiply this by 50 to get your answer

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

first multiply 20 by 50 then divide by 100

OpenStudy (anonymous):

It would be 10 tickets because 10/50=20/100 You would divide 100 by 2 to get 50 and what you do to the bottom, you do to the top. So you divide 20 by 2 to get 10 Therefore, ten tickets